One end of a line segment is (1,-2). If the midpoint of the segment is (4,3), find the coordinate of the other endpoint of the segment.

Thanks for the big help.. Kindly put the solution..

Answer by stanbon(75887)

One end of a line segment is (1,-2). If the midpoint of the segment is (4,3), find the coordinate of the other endpoint of the segment.
---------------------
The coordinates of the mid-point are the average of the coordinates
of the endpoints.
----------------------------
Let the other endpoint be (x,y)
Equations:
(x+1)/2 = 4
x+1 = 8
x = 7
-----------------
(y+-2)/2 = 3
y-2 = 6
y = 8
-------------
The other endpoint is (7,8)
=================================
Cheers,
Stan H.